Shift report 20010808 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed Aug 08 2001 - 16:27:09 EDT

  • Next message: Djamel Ouerdane: "Calibration"

    ------------------------------------------------------------
    Date: Wed Aug  8 16:27:08 2001, Shift: 08:00-16:00
    Supervisor: JHL, with: 
    Log book#7 pages 47 - 48
    
    Shift Summary:
    ------------------------------------------------------------
    1. Machine/Beam
       - Cogged Beam until 9:40
       - Since then no useful beam 
       - Machine is off due to power-saving  
         for 14:00 - 18:00 (Another HOT day here!)
    
    2. Data
       - Physics Data (including from the previous shift)
         for Runs 4064 - 4069 (~400K triggers)
    
    3. Detector Problems/Fix
       - Another broken wire found in T1
         and has been removed (RD,FV)
       - T1 cabling is being modified to run with voltages 
         on Field shaping wires (RD,FV)
       - Bob fixed the T2 read-out problem
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Wed Aug 08 2001 - 16:27:51 EDT